# supplement

### v1

@@ 1 Introduction
where $\mathbb{S}$ denotes a simplex, which is a $(D-1)$-dimensional subset of real-valued space $R^{D-1}$, denotes vector transpose, and $c$ is a fixed constant, set typically to one (portions) (Scealy et al., 2015), $10^{6}$ parts per million (Scealy et al., 2015), or $10^{5}$ life-table death counts.
Due to the non-negativity and summability constraints inherent in compositional data, conventional linear techniques may be inappropriate, as they do not account for the unique variance-covariance structure of such data. A common strategy is to first transform the data to remove the summability constraint, thereby enabling the application of linear methods. One widely used approach is the family of log-ratio transformations, particularly the centered log-ratio (CLR) transformation (see, e.g., Aitchison, 1986). An alternative is the cumulative distribution function (CDF) transformation, which leverages monotonicity to effectively handle the presence of zero values (see Shang & Haberman, 2025).
This paper makes two key contributions. First, we evaluate and compare the point forecast accuracy of the CLR and CDF transformations, using data from 24 countries in the Human Mortality Database (2025), with time series starting in or before 1950. Second, while the existing literature has largely focused on point forecasts, we introduce a computationally efficient approach for constructing pointwise prediction intervals and compare the interval forecast accuracy between the CLR and CDF transformations.
@@ 3.1 Centered log-ratio transformation
    $\alpha^{s}_{n,x}=\exp\Big{(}\frac{1}{n}\sum^{n}_{t=1}\ln d^{s}_{t,x}\Big{)}.$
    We compute the centered data by dividing each death count by its geometric mean
@@
    $\beta_{t,x}^{s}=\sum^{L_{s}}_{\ell=1}\gamma_{t,\ell}\phi_{\ell,x}^{s}+\omega_{t,x}^{s},$
    where $\phi_{\ell,x}^{s}$ is the $\ell$th estimated principal component for age $x$, $\gamma^{s}_{t,\ell}=\langle\beta_{t,x}^{s},\phi_{\ell,x}^{s}\rangle$ is the $\ell$th estimated principal component score at time $t$, $\omega_{t,x}^{s}$ denotes the model residuals for age $x$ in year $t$, and $L_{s}$ denotes the number of retained principal components. For selecting the number of components $L_{s}$, we implement an eigenvalue ratio criterion of Li et al. (2020), as well as $L_{s}=6$ recommended in Hyndman et al. (2013).
@@
    where $\widehat{\gamma}_{n+h|n,\ell}$ denotes the $h$-step-ahead forecast of the principal component scores.
@@ 3.2 Cumulative distribution function transformation
    $Z^{s}_{t,x}=\text{logit}(D^{s}_{t,x})=\ln\left(\frac{D^{s}_{t,x}}{1-D^{s}_{t,x}}\right),$
    where $\ln(\cdot)$ denotes the natural logarithm.
@@
    By conditioning on the estimated functional principal components $\bm{\Psi}^{s}_{x}=(\psi_{1,x},\dots,\psi_{K_{s},x})$ and observed data $\bm{Z}_{x}^{s}=(Z_{1,x}^{s},\dots,Z_{n,x}^{s})$, the $h$-step-ahead forecast of $Z_{n+h,x}^{s}$ can be obtained
@@ 4.2 Point forecast evaluation metrics
$\text{KLD}(h)=$ $D_{\text{KL}}(d_{n+\xi,x}||\widehat{d}_{n+\xi|n,x})+D_{\text{KL}}(\widehat{d}_{n+\xi|n,x}||d_{n+\xi,x})$
=\ = $\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}d_{n+\xi,x}\cdot(\ln d_{n+\xi,x}-\ln\widehat{d}_{n+\xi|n,x})+$
@@ 4.4 Construction of pointwise prediction intervals
Using the data in the training sample, we carry out an expanding window approach to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,\dots,19$. For different forecast horizons, we have different numbers of curves in the validation set. For $h=1$, we have 20 years in evaluate the forecast errors; when $h=19$, we have two years to evaluate the residual functions between the holdout data in the validation set and their forecasts. Based on these residuals, we can compute the functional standard deviation, which requires at least two years of data. The forecast errors can be denoted by $\widehat{\epsilon}_{\nu,x}=d_{\nu,x}-\widehat{d}_{\nu,x}$, where $\widehat{d}_{\nu,x}$ denotes the forecasts obtained from the two transformations, and $\nu=1,\dots,(21-h)$ denotes the number of observations in the validation set for $h=1,\dots,19$.
Let us denote $\Gamma_{x}=\text{sd}(\widehat{\epsilon}_{\nu,x})$, where $\text{sd}(\cdot)$ can be pointwise standard deviation. We denote $\bm{\Gamma}=(\Gamma_{1},\dots,\Gamma_{111})^{\top}$ and $\bm{\widehat{\epsilon}}_{\nu}=(\widehat{\epsilon}_{\nu,1},\dots,\widehat{\epsilon}_{\nu,111})^{\top}$. For a level of significance $\alpha$, customarily $\alpha=0.2$ or 0.05, we aim to find $\theta_{\alpha}$ such that approximately $100(1-\alpha)\%$ of the residuals satisfy
$-\theta_{\alpha}\bm{\Gamma}\leq\bm{\widehat{\epsilon}}_{\nu}\leq\theta_{\alpha}\bm{\Gamma},\qquad v=1,\dots,(21-h).$
@@
$\text{Pr}(-\theta_{\alpha}\Gamma_{x}\leq d_{n+h,x}-\widehat{d}_{n+h|n,x}\leq\theta_{\alpha}\Gamma_{x})\approx\frac{1}{21-h}\sum^{21-h}_{\nu=1}\mathds{1}(-\theta_{\alpha}\bm{\Gamma}\leq\bm{\widehat{\epsilon}}_{\nu}\leq\theta_{\alpha}\bm{\Gamma}),$
@@
To determine the optimal $\theta_{\alpha}$, the samples in the validation set are used to calibrate a prediction interval so that its empirical coverage probability is close to its nominal coverage probability. With the estimated $\theta_{\alpha}$, the $100(1-\alpha)\%$ prediction interval of $d_{n+h,x}$ can be obtained as
$\big{[}\widehat{d}_{n+h|n,x}-\theta_{\alpha}\Gamma_{x},\quad\widehat{d}_{n+h|n,x}+\theta_{\alpha}\Gamma_{x}\big{]}.$
@@ 4.5 Interval forecast evaluation metrics
where $S_{\alpha}(\widehat{d}_{n+\xi|n,x}^{\text{lb}},\widehat{d}_{n+\xi|n,x}^{\text{ub}},d_{n+\xi,x})$ denotes the interval score at the $\xi$th curve in the forecasting period. Since we require at least two curves to compute standard deviation, the forecast horizon is $h=1,2,\dots,19$.
@@ 4.6 Comparison of interval forecast accuracy
While Table LABEL:tab:2 compares the interval forecast accuracy based on the CPD and mean interval scores, aggregated over 24 countries. In Figure 4, we provide several heatmaps that show the number of times one method produces the smallest errors. For modeling the female data, the difference between the CLR and CDF transformations is marginal. For modeling the male data, the CDF transformation with $K=6$ is advocated with the smallest CPD and mean interval score.
@@ 5 Comparisons of life expectancy at birth and annuity price
With the CDF transformation, we obtain 50-years-ahead forecasts of life-table death counts for all countries. As a vehicle of demonstration, we display the forecasts for Australia in Figure 5 and observe a continuing trend of further longevity. These forecasts are time-series extrapolations of the past trend, and thus should be used with caution. Any extreme shocks, such as COVID-19, are likely to have a substantial impact on the mortality trend.
@@
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(d_{x})$ (from 2022 to 2071) to their corresponding survival probabilities ($p_{x}$). We plug the age-specific survival probabilities into the calculation of single-premium term immediate annuities, and we adopt a cohort approach to calculating the survival probabilities, as is the practice of insurance companies, that is, survival probability $p_{60}$ is computed from the forecasted life-table death counts in 2022, $p_{61}$ is computed from the forecasted life-table death counts in 2023, $\dots$, $p_{109}$ is computed from the forecasted life-table death counts in 2071. The $\tau$ year survival probability of a person with entry age $x$ at the contract time is determined by
@@
where $d_{x+j-1}$ denotes the number of death counts between ages $x+j-1$ and $x+j$, and $l_{x+j-1}$ represents the number of lives alive at age $x+j-1$.
@@
Figure 8 presents dotcharts for the estimated annuity prices of the 24 countries at entry age 60 with a maturity period of 5 years. With a return of one monetary amount, the estimated annuity price is the highest for Japan, while the two with the lowest values are Iceland and Hungary.
@@
Figures 10a and 10b show the 80% and 95% prediction intervals for annuity prices at entry age 60 with 5-year maturity across 24 countries, while Figures 10c and 10d show the results for entry age 70 with 20-year maturity. With a return of one monetary unit, estimates are shown as open circles, with open intervals indicating prediction intervals. Forecast uncertainty is noticeably higher for Iceland and lower for Hungary. With increasing maturity, prediction intervals widen, reflecting greater long-term uncertainty.
@@ 6 Conclusion
There are a few ways in which the methodology presented in the paper could be extended, and we briefly discuss two. First, although we calculate temporary annuities, other types of annuity prices, such as the whole-life immediate annuity or deferred annuity, are possible. Second, we could consider cohort life-table death counts for modeling a group of individuals, such as baby boomers.
Acknolwedgment

### v2

# Forecasting age distribution of deaths across countries: Life expectancy and annuity valuation

@@ 1 Introduction
where $\mathbb{S}$ denotes a simplex, which is a (D1)(D-1)-dimensional subset of real-valued space $R^{D-1}$, denotes vector transpose, and $c$ is a fixed constant, set typically to one (portions) (Scealy et al., 2015), $10^{6}$ parts per million (Scealy et al., 2015), or $10^{5}$ life-table death counts.
Due to the non-negativity and summability constraints inherent in compositional data, conventional linear techniques may be inappropriate, as they do not account for the unique variance-covariance structure of such data. A common strategy is to first transform the data to remove the summability constraint, thereby enabling the application of linear methods. One widely used approach is the family of log-ratio transformations, particularly the centered log-ratio (CLR) transformation (see, e.g., Aitchison, 1986). An alternative is the cumulative distribution function (CDF) transformation, which leverages monotonicity to effectively handle the presence of zero values (see Shang & Haberman, 2025b).
This paper makes two key contributions. First, we evaluate and compare the point forecast accuracy of the CLR and CDF transformations, using data from 24 countries in the Human Mortality Database (2025), with time series starting in or before 1950. Second, while the existing literature has largely focused on point forecasts, we implement a computationally efficient approach of Shang & Haberman (2025a) for constructing pointwise prediction intervals and compare the interval forecast accuracy between the CLR and CDF transformations. From a forecasting perspective, this study presents a comprehensive comparison across countries and further explores heterogeneity in mortality and its implications for life expectancy forecasting and annuity pricing.
@@ 3.1 Centered log-ratio transformation
    $\alpha^{s}_{n,x}=\exp\Big{(}\frac{1}{n}\sum^{n}_{t=1}\ln d^{s}_{t,x}\Big{)},$
    where $\ln(\cdot)$ denotes the natural logarithm. We compute the centered data by dividing each death count by its geometric mean
@@
    $\beta_{t,x}^{s}=\sum^{L_{s}}_{\ell=1}\gamma^{s}_{t,\ell}\phi_{\ell,x}^{s}+\omega_{t,x}^{s},$
[…]
@@
    where $\widehat{\gamma}_{n+h|n,\ell}$ denotes the $h$-step-ahead forecast of the principal component scores. These forecasts can be obtained via a univariate time-series forecasting method, such as the exponential smoothing method of Hyndman et al. (2008). The automatic algorithm of Hyndman et al. (2002) can be used to select the optimal exponential smoothing model among 15 possible candidates (see Table 1 Hyndman & Khandakar, 2008), based on the corrected Akaike information criterion.
@@ 3.2 Cumulative distribution function transformation
    $Z^{s}_{t,x}=\text{logit}(D^{s}_{t,x})=\ln\left(\frac{D^{s}_{t,x}}{1-D^{s}_{t,x}}\right).$
@@
    By conditioning on the estimated functional principal components $\bm{\Psi}^{s}_{x}=(\psi^{s}_{1,x},\dots,\psi^{s}_{K_{s},x})$ and observed data $\bm{Z}_{x}^{s}=(Z_{1,x}^{s},\dots,Z_{n,x}^{s})$, the $h$-step-ahead forecast of $Z_{n+h,x}^{s}$ can be obtained
@@ 4.2 Point forecast evaluation metrics
$\text{KLD}(h)=\$$D_{\text{KL}}(d_{n+\xi,x}||\widehat{d}_{n+\xi|n,x})+D_{\text{KL}}(\widehat{d}_{n+\xi|n,x}||d_{n+\xi,x})$
$=\$$\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}d_{n+\xi,x}\cdot(\ln d_{n+\xi,x}-\ln\widehat{d}_{n+\xi|n,x})+$
@@ 4.3 Comparison of point forecast accuracy
In Figure 3, we demonstrate the forecasts of the life-table death counts obtained from the CLR and CDF transformations. For one-step-ahead forecasts, the differences between the two transformation are marginal. As forecast horizon increases to $h=20$, the difference is more visible. As measured by the KLD and JSD, the CDF transformation produces smaller forecast errors than those from the CLR transformation for $h=1$ and $h=20$ considered.
Figure 3: One-step-ahead and 20-step-ahead forecasts of life-table death counts for Australia women (first row) and men (second row).
@@ 4.4 Construction of pointwise prediction intervals
Using the data in the training sample, we carry out an expanding window approach to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,\dots,19$. For different forecast horizons, we have different numbers of curves in the validation set. For $h=1$, we have 20 years in evaluate the forecast errors; when $h=19$, we have two years to evaluate the residual functions between the holdout data in the validation set and their forecasts. Based on these residuals, we can compute the functional standard deviation, which requires at least two years of data. The forecast errors can be denoted by $\widehat{\epsilon}_{\nu,x}=d_{\nu,x}-\widehat{d}_{\nu,x}$, where $\widehat{d}_{\nu,x}$ denotes the forecasts obtained from the two transformations, and $\nu=1,\dots,(21-h)$ denotes the number of observations in the validation set for $h=1,\dots,19$. When $h=1$, we have 20 observations in the validation set; when $h=19$, we have two observations.
Let us denote $\Gamma_{\nu,x}=\text{sd}(\widehat{\epsilon}_{\nu,x})$, where $\text{sd}(\cdot)$ can be pointwise standard deviation. We denote $\bm{\Gamma}_{\nu}=(\Gamma_{\nu,1},\dots,\Gamma_{\nu,111})^{\top}$ and $\bm{\widehat{\epsilon}}_{\nu}=(\widehat{\epsilon}_{\nu,1},\dots,\widehat{\epsilon}_{\nu,111})^{\top}$. For a level of significance $\alpha$, customarily $\alpha=0.2$ or 0.05, we aim to find $\theta_{h,\alpha}$ such that approximately $100(1-\alpha)\%$ of the residuals satisfy
$\theta_{h,\alpha}\bm{\Gamma}_{\nu}\leq\bm{\widehat{\epsilon}}_{\nu}\leq\theta_{h,\alpha}\bm{\Gamma}_{\nu},\qquad v=1,\dots,(21-h).$
@@
$\text{Pr}(-\theta_{h,\alpha}\Gamma_{\nu,x}\leq d_{n+h,x}-\widehat{d}_{n+h|n,x}\leq\theta_{h,\alpha}\Gamma_{\nu,x})\approx\frac{1}{(21-h)\times 111}\sum^{21-h}_{\nu=1}\sum^{111}_{x=1}\mathds{1}(-\theta_{h,\alpha}\Gamma_{\nu,x}\leq\widehat{\epsilon}_{\nu,x}\leq\theta_{h,\alpha}\Gamma_{\nu,x}),$
@@
To determine the optimal $\theta_{h,\alpha}$, the samples in the validation set are used to calibrate a prediction interval so that its empirical coverage probability is close to its nominal coverage probability. With the estimated $\theta_{h,\alpha}$, the $100(1-\alpha)\%$ prediction interval of $d_{n+h,x}$ can be obtained as
$\big{[}\widehat{d}_{n+h|n,x}-\theta_{h,\alpha}\Gamma_{\nu,x},\quad\widehat{d}_{n+h|n,x}+\theta_{h,\alpha}\Gamma_{\nu,x}\big{]},$
where $v$ denotes the number of observations in the validation set. As forecast horizon $h$ increases, the value of $\theta_{h,\alpha}$ generally increases.
@@ 4.5 Interval forecast evaluation metrics
[…]
@@ 4.6 Comparison of interval forecast accuracy
While Table LABEL:tab:2 compares the interval forecast accuracy based on the CPD and mean interval scores, aggregated over 24 countries. Based on the averaged CPDα and $\overline{S}_{\alpha}$ across various forecast horizons, the CDF transformation provides the smaller interval forecast errors than those obtained from the CLR transformation. In Figure 5, we provide several heatmaps that show the number of times one method produces the smallest errors. For modeling the female data, the difference between the CLR and CDF transformations is marginal. Across forecast horizons, the inconsistency results between the two transformation may be because of the optimal estimation of $\theta_{h,\alpha}$ parameters in both transformations. For modeling the male data, the CDF transformation with $K=6$ is advocated with the smallest CPD and mean interval score.
@@ 5 Comparisons of life expectancy at birth and annuity price
Since the CDF transformation generally provides smaller forecast errors than those obtained from the CLR transformation. Using the CDF transformation, we obtain 50-years-ahead forecasts of life-table death counts for all countries. As a vehicle of demonstration, we display the forecasts for Australia in Figure 6 and observe a continuing trend of further longevity. These forecasts are time-series extrapolations of the past trend, and thus should be used with caution. Any extreme shocks, such as COVID-19, are likely to have a substantial impact on the mortality trend.
@@
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(\bm{d}_{x})$ to their corresponding survival probabilities ($\bm{p}_{x}$) from 2022 to 2071. We plug the age-specific survival probabilities into the calculation of single-premium term immediate annuities, and we adopt a cohort approach to calculating the survival probabilities, as is the practice of insurance companies, that is, survival probability $p_{60}$ is computed from the forecasted life-table death counts in 2022, $p_{61}$ is computed from the forecasted life-table death counts in 2023, $\dots$, $p_{109}$ is computed from the forecasted life-table death counts in 2071. The $\tau$ year survival probability of a person with entry age $x$ at the contract time is determined by
@@
where $d_{x+j-1}$ denotes the number of death counts between ages x+j1x+j-1 and $x+j$, and $l_{x+j-1}$ represents the number of lives alive at age x+j1x+j-1.
@@
Figure 9 presents dotcharts for the estimated annuity prices of the 24 countries at entry age 60 with a maturity period of 5 years. With a return of one monetary amount, the estimated annuity price is the highest for Japan, while the two countries with the lowest values are Iceland and Hungary.
@@
Figures 11a and 11b show the 80% and 95% prediction intervals for annuity prices at entry age 60 with 5-year maturity across 24 countries, while Figures 11c and 11d show the results for entry age 70 with 20-year maturity. With a return of one monetary unit, estimates are shown as open circles, with open intervals indicating prediction intervals. Forecast uncertainty is noticeably higher for Iceland and lower for Hungary. With increasing maturity, prediction intervals widen, reflecting greater long-term uncertainty.
@@ 6 Conclusion
There are a few ways in which the methodology presented in the paper could be extended, and we briefly discuss four. 1) When forecasting principal component scores, we implement exponential smoothing. Other univariate or multivariate time-series forecasting methods may also be plausible. 2) We consider two transformations, namely the centered log-ratio and cumulative distribution function transformations. There exist other transformations that could also be considered, namely additive or isometric log-ratio (Greenacre et al., 2023). While these transformations are variants of the centered log-ratio, we could also consider intrinsic statistical models, such as Wasserstein autoregression (Zhang et al., 2022) or optimal transport (Zhu & Müller, 2023), for modeling density-valued time series. 3) Although we calculate temporary annuities, other types of annuity prices, such as the whole-life or deferred annuity, are possible. 4) We could consider cohort life-table death counts for modeling a particular group of individuals, such as baby boomers.
Data availability statement
The primary source for the analysis is the freely available data in the Human Mortality Database (HMD) that is publicly accessible subject to creating a free HMD account. No further pre-processing was applied to this data. For reproducibility, the computer [Uncaptioned image] code is available at https://github.com/hanshang/CLR_vs_CDF_transformation.
Funding statement